LASER ABSORPTION SPECTROSCOPY OF HYDROCARBON FLAMES

\underline{SERGEY CHESKIS}, IGOR DERZY, School of Chemistry, Tel Aviv University, Tel Aviv 69978, Israel; VLADIMIR A. LOZOVSKY, Semenov Institute of Chemical Phisics, Russian Academy of Sciences, 4 Kosygina str., 117977 Moscow, Russia.

Intracavity Laser Absorption Spectroscopy (ICLAS) and Cavity Ring-Down Spectroscopy (CRDS) were used to detect absorption spectra of CH(C^{2}Σ^{+} ← X^{2}Π) at 314 nm, ^{1}CH_{2} (b̃^{1}B_{1} ← ã^{1}A_{1}) at 590 and 620 nm, NH (A^{3}Π_{i} ← X^{3}Σ^{−}) at 336 nm , and NH_{2} (Ã^{2}A_{1} ← X̃^{2}B_{1}) at 598 nm in a low-pressure (30 Torr) stoichiometric methane/oxygen/nitrogen flat flame doped with a small amount of nitrous oxide . The CH and NH radicals were monitored by CRDS whereas 1CH_{2} and NH_{2} were monitored by ICLAS. The absolute concentration profiles of those radicals were measured. The radical absorption spectra were recorded with good signal-to-noise ratio. The spectra of the ^{1}CH_{2} radical were measured in different spectral ranges that allowed us better determination of its absorption cross section. For the first time the absolute concentrations of NH and NH_{2} were measured in the flames of this kind. The agreement between experimental results and model predictions based on the GRI-Mech 2.11 mechanism is discussed.